# SoK: Membership Inference Attacks on LLMs are Rushing Nowhere (and How to Fix It)

### v2

@@ 3 Literature review
In Table 1, we review prior work studying MIAs against LLMs. We focus on the most recent work (’23-’24) and also include initial seminal work.
@@
First, we distinguish between MIAs against finetuned and pretrained models, as defined in Sec. 2. We find that most recent MIAs are primarily targeting pretrained LLMs rather than models finetuned on smaller datasets.
@@
Second, we distinguish between MIAs based on the unit of interest they consider. Most commonly, prior work has considered sequence-level MIAs as defined in Sec. 2. Some exceptions include scaling MIAs to the level of a document [86, 69] or dataset [67]. In the rest of this work we focus primarily on sequence-level MIAs, consistent with the majority of the literature, and will revisit document-level MIAs below.
@@
Third, we report all studies developing new MIAs; i.e. designing novel scoring functions $f_{\text{seq}}(\mathcal{M},s_{t})$, which would more effectively capture the likelihood of membership [16, 106, 68, 86, 109, 109, 101, 62, 105, 108, 97, 73, 45, 55]. Studies not proposing new MIAs still evaluate MIAs on particular setups, e.g. to evaluate the effectiveness of privacy-preserving technologies [17, 79], detect the use of copyrighted content [70, 100] or draw connections between MIAs and PII leakage [65]. In Table 1 we show that there is a significant body of recent work focused on developing new MIAs.
@@
Fourth, we distinguish the primary motivation of each MIA study as either privacy and security, copyright or dataset contamination. While initial work was primarily studying MIAs from a privacy and security perspective [15, 16, 71, 68, 65], we find that recent work increasingly considers novel questions associated with LLM pretraining, i.e. detecting the use of copyright-protected content or evaluation benchmarks.
@@
Finally, across all works, we also classify the datasets which have been used to study MIAs. Importantly, we identify that most of the recent studies have relied on datasets collected post-hoc for MIA evaluation ().
4 Current evaluation methods are flawed
@@
In this section, we examine the validity of using datasets collected post-hoc for MIA evaluation, as employed by most recent work ( in Table 1).
4.1 MIAs relying on post-hoc data collection
@@
Recent work [86, 69] has developed strategies to create a set of likely members and likely non-members using information about training data sources and the LLM release date. We refer to such data collection strategies as post-hoc, where an attacker aims to infer membership for a released LLM, without access to the randomized training data distribution. In both works, members are collected by sampling documents from the often publicly available datasets widely used for LLM training (e.g. Wikipedia articles [86], books in the public domain from Project Gutenberg [^3] https://www.gutenberg.org/, ArXiv papers [^4] https://arxiv.org/ [69]). Non-members are collected as documents from the same data sources but released after the LLM’s training data cutoff date. Using this collection technique for evaluation, Shi et al. [86] propose a novel MIA scoring function based on the lowest predicted probabilities returned by the LLM. Authors achieved a membership AUC of $0.74$ for model LLaMA-30B [94] on the WikiMIA dataset, and an AUC of $0.88$ for GPT-3 [10] on copyright-protected books (BookMIA). Meeus et al. [69] train and evaluate a custom membership classifier against OpenLLaMA [32] and achieved MIA AUC of $0.86$ and $0.68$ for books and academic papers, respectively.
@@
Such post-hoc approaches to collect datasets for MIA evaluation are not strictly randomized, raising the question whether they sufficiently approximate randomness in a privacy game or if biases may be introduced during the data collection process. Indeed, doubts have recently been raised on whether this construction of members and non-members constitutes a distribution shift due to the temporal gap in data collection [27, 67]. Even if efforts to mitigate any distribution shift are employed (e.g. Meeus et al. [69] ensured that the original publication dates of member and non-member books were similarly distributed), post-hoc data collection leaves the possibility of a distribution shift. If this is the case, the membership classifier might learn the temporal gap in language rather than membership, potentially questioning the validity of MIA results obtained.
4.2 Quantifying a potential distribution shift
@@
Duan et al. [27] was the first to flag that the previously reported high MIA AUC might be due to the distribution shift between members and non-members. They observe a significant change in $7$-gram overlap between members and non-members, and also show how the MIA AUC increases when the temporal gap between members and non-members increases. Shortly afterwards, Maini et al. [67] also flags that post-hoc data collection introduces an important confounder, making members and non-members not IID.
@@
We here propose to train and evaluate a bag of words classifier to distinguish members from non-members. We argue that such classifier can be used to measure what performance for this binary prediction task can be reached purely based on the textual data without taking the target model into account. If such baseline is better than a random guess, this would be indicative of a shift in language.
@@
We use a random forest classifier ($500$ trees with maximum depth of $2$ and minimum samples per leaf of $10$) with the count of words ($1$-grams) appearing in at least $5\%$ of the training documents as features. For each dataset, we train the classifier on $80\%$ of all documents, evaluating it on the remaining $20\%$ (averaged over $5$ runs, with balanced membership).
@@
We instantiate the bag of words classifier on multiple datasets collected post-hoc and used to evaluate MIAs against LLMs. We consider all (available) datasets: the WikiMIA and BookMIA datasets as introduced by Shi et al. [86], the books from Project Gutenberg and academic papers from ArXiv used by Meeus et al. [69], ArXiv papers from ArXivMIA [62] and StackExchange samples from StackMIA [105].
Dataset AUC Predictive words (sample)
@@
Table 2: AUC (mean±$\pm$±standard deviation) for the bag of words classifier across datasets collected post-hoc.
Table 2 shows that the bag of words classifier can distinguish members from non-members significantly better than a random guess (AUC of $0.5$). For some datasets, the classifier reaches near-perfect AUC, i.e. $0.99$ for the WikiMIA dataset, $0.94$ for the BookMIA dataset and $0.97$ for the books from Project Gutenberg (full), strikingly exceeding the MIA AUC reported by attacks utilizing the corresponding target model [86, 69]. For other datasets, the performance is relatively lower, yet remains significantly better (AUC of $\geq 0.6$) than a random guess baseline. This means that, all datasets relying on post-hoc data collection which have been released, suffer from a significant and often strong distribution shift. Moreover, the simplicity of the bag of words classifier suggests that it likely provides a lower bound on classification performance. More advanced classifiers could potentially detect other, subtler distribution shifts.
@@
For WikiMIA [86], most predictive words correspond to years. As the dataset is constructed using text from Wikipedia ‘events‘, entries almost always contain dates. Data collected for different time periods can thus be perfectly distinguished by the bag of words classifier using primarily the years. Similarly for ArXiv papers used by Meeus et al. [69], years (in citations) carry the most predictive value. For ArXivMIA [62], the bias in years gets eliminated (they use abstracts, which rarely contain dates). Yet, we recover an AUC of $0.60$ with the most predictive words suggesting a change in topics over time (increased use of ’training’ and ’github’ in recent work).
@@
Books, on the other hand, likely contain less dates. Meeus et al. [69] even attempted to control for a potential distribution shift by ensuring publication year to be similarly distributed between 1850 and 1910 for members and non-members. Yet, in this case, the AUC is near-perfect with the most predictive words relating to template text or formatting. We find that the beginning and end of the books from Project Gutenberg include certain templates (e.g. ‘This ebook was transcribed‘), which have changed over time, allowing the classifier to distinguish members from non-members. Similarly for BookMIA [86], certain formatting styles have changed between members and non-members (e.g. ‘did not‘ rather than ‘didn’t‘).
@@
The distribution shift can be very subtle. We attempt to mitigate the distribution shift due to the template text by excluding the first and last 10% of the books (mid). Table 2 shows, however, that the AUC of the bag of words classifier does decrease from $0.97$, yet remains as high as $0.89$. When it comes to the most predictive words, we find that books, while originally published in the same time period but released online at a later stage, contain less old formatting and notably, less politically incorrect words such as the n-word.
@@
Our results show that a simple bag-of-words classifier can consistently distinguish between members and non-members across all observed datasets. The analysis of the predictive words reveals that these biases can be subtle, and hard to mitigate. This implies that a fundamental assumption in membership inference, i.e. the members and non-members are IID, does not hold. The post-hoc collection of members and non-members thus fails to provide a proper, randomized privacy game to evaluate MIAs. We are hence unable to exclusively attribute the high MIA AUC reported by studies relying on post-hoc data collection [69, 86, 109, 101, 62, 105, 108, 97, 73, 45] to memorization of the LLM.
4.3 Perpetuating flaws in subsequent research
@@
Since the initial work relying on data collected post-hoc [86, 69], we identify at least 10 novel attack methodologies evaluated using datasets collected in a post-hoc manner [83, 109, 101, 62, 28, 105, 108, 45, 97], the latest of which proposed (shortly before submission) in September ’24 [97]. This work either uses datasets released by Shi et al. [86] or newly created datasets of members and non-members in a similar, post-hoc manner: ArXivMIA [62], StackMIA [105], Dolma-Book [108] and Wikipedia Cutoff [73].
@@
With the results of our bag of words classifier (Table 2), we find that all (accessible) datasets collected post-hoc suffer from distribution shift between members and non-members, which renders any comparison of MIA methodologies unreliable. Indeed, if one MIA method outperforms the other, it is impossible to determine whether one MIA is better at distinguishing which data points are memorized by the target LLM or if it is simply better at identifying the distribution shift.
@@
The problem also extends beyond the development of new MIAs, with numerous subsequent work using similar datasets for evaluating the effectiveness of machine unlearning [103] or watermarking techniques [48, 100] as a privacy-preserving technology [79]. The use of this flawed MIA evaluation setup in broader contexts is equally concerning. It may lead to false conclusions about the effectiveness of privacy-preserving technologies, which could have far-reaching implications for LLM development and deployment.
5 Randomization for MIAs against LLMs
@@
5.1 Why is it different than for MIAs against ML models?
@@
MIAs have been extensively applied to various domains of ML, including image classification [88, 14, 84, 96, 60] and natural language processing [85, 36, 93, 89], all evaluated in randomized privacy games. Ye et al. [104] distinguishes between a membership inference game involving a fixed record and one involving a fixed model. In the fixed record scenario, multiple ML models are trained with the target record either included or excluded, commonly referred to as shadow modeling [63, 64, 91, 41, 12, 82, 106]. For instance, Shokri et al.[88] used up to $100$ shadow models for image classification models, while Song et al.[90] trained 10 shadow models to evaluate privacy in smaller language models (RNNs). However, replicating this approach for LLMs is impractical due to the extremely high costs of pretraining. For example, training Llama-3.1-405B on 15 trillion tokens required 30.84 million GPU hours  [3]. Such resource demands make privacy games relying on multiple models infeasible.
@@
As a result, privacy games to evaluate MIAs against LLMs likely rely on the fixed model scenario, in which a single ML model is trained with some randomly sampled target records from the same distribution, either included or excluded. While this setup has been widely examined in previous MIAs against ML [88, 18, 76, 84, 98], their practical implementation for LLMs is not without its limitations, which we will address below.
5.2 Unit of interest to infer membership
@@
In traditional MIAs against ML models, the unit of interest to infer membership is naturally defined (e.g. single record, image). For text, most prior work considered sequences of a fixed, often truncated length as the unit of interest. For instance, Carlini et al. [16] use sequences of $256$ tokens, while Shi et al. [86] and subsequent work [109, 101, 62, 105, 108, 97, 73, 45] consider random excerpts between $32$ and $256$ tokens. When injecting randomized sequences, prior work has considered sequences up to $80$ [100] or $100$ [70] tokens.
@@
In some cases, however, the motivation of applying MIAs against LLMs comes down to inferring membership of larger pieces of text, from the level of a document [69, 86] to an entire dataset [67]. For instance, in the context of copyright, the task of interest is inferring whether the entire content (e.g. news articles, book) was used rather than certain excerpts. This has motivated prior work [69, 86, 67] to instantiate document- or dataset-level membership inference, aiming to meaningfully aggregate membership signal to larger pieces of text.
@@
A document-level MIA aims to determine whether a given document $d_{t}$ has been used to train target LLM $\mathcal{M}$. In Sec. 2, we formalize a document as an ordered sequence of tokens $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, of length $N_{d_{i}}$, which originates from a single, identifiable source and is coherent and semantically meaningful when considered in isolation. Given an LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$ and a target document $d_{t}$, a document-level MIA is a scoring function $f_{\text{doc}}(\mathcal{M},d_{t})$ such that $\hat{y}_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }f_{\text{doc}}(\mathcal{M},d_{t})\geq\tau_{\text{doc}}\text{, and }0\text{ otherwise}$. The true membership label is $y_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }d_{i}\in\mathcal{D}\text{, and }0\text{ otherwise}$. For simplicity, we argue that the task of dataset inference as proposed by Maini et al. [67] is a generalization of document-level MIAs.
@@
Document-level MIAs differ from sequence-level MIAs against LLMs, as aggregating meaningful signal across significantly more tokens might be challenging. Multiple ways have been proposed to compute $f_{\text{doc}}(\mathcal{M},d_{t})$. Meeus et al. [69] iterates through the entire document $d_{t}$, queries the LLM on truncated chunks, and aggregates all token-level probabilities to multiple document-level features as input for a membership classifier. Shi et al. [86] proposes to sample $k$ random excerpts, i.e. sequences $s_{tj}$, from document $d_{t}$ to compute the sequence-level scoring function $f_{\text{seq}}(\mathcal{M},s_{tj})$ for all $j=1,\ldots k$. The document-level scoring function is computed as an average binary prediction $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$, where $\tau_{\text{seq}}$ is determined to maximize accuracy on the sequence-level MIA. Lastly, Maini et al. [67] aggregates values of the sequence-level scoring function to a statistical test of the document level as $y_{\text{doc}}(\mathcal{M},d_{t})$.
5.3 Overlap between members and non-members

### v3

@@ III Literature review
In Table I, we review the literature studying MIAs against LLMs. We focus on the most recent work (’23-’24) while also including initial seminal work.
@@
First, we distinguish between MIAs applied to finetuned and pretrained models, as defined in Sec. II. We find that most recent MIAs are primarily targeting pretrained LLMs rather than models finetuned on smaller datasets.
@@
Second, we distinguish between MIAs based on the unit of interest they consider. Most commonly, prior work has considered sequence-level MIAs, while some exceptions include scaling MIAs to the level of a document [85, 68] or dataset [65]. In the rest of this work we focus primarily on sequence-level MIAs, consistent with the majority of the literature, and will revisit document-level MIAs below (Sec. V-B).
@@
Third, we distinguish between studies developing new MIA methodologies (i.e. designing novel scoring functions $f_{\text{seq}}(\mathcal{M},s_{t})$) and studies merely evaluating prior MIA methodologies in specific setups. In Table I we show that there is a significant body of recent work focused on developing new MIAs, often claiming to more effectively capture the likelihood of membership than prior work [15, 105, 66, 85, 109, 109, 100, 60, 104, 107, 96, 72, 45, 54]. Studies not proposing new MIAs still evaluate MIAs on particular setups, e.g. to evaluate the effectiveness of privacy-preserving technologies [17, 78], detect the use of copyrighted content [69, 99] or draw connections between MIAs and PII leakage [63].
@@
Fourth, we distinguish the primary motivation of each MIA study as either privacy and security, copyright or dataset contamination. While initial work was primarily studying MIAs from a privacy and security perspective [14, 15, 70, 66, 63], we find that recent work increasingly considers novel questions associated with LLM pretraining, for instance detecting the inclusion of copyright-protected content or evaluation benchmarks.
@@
Finally, we also classify the datasets which have been used to study MIAs. Importantly, we identify that most of the recent studies have relied on datasets collected post-hoc for MIA evaluation ().
IV Current evaluation methods are flawed
@@
In this section, we introduce a bag of words classifier as a model-less baseline to distinguish members from non-members and use this to examine the validity of using datasets collected post-hoc for MIA evaluation, as employed by most recent work ( in Table I).
IV-A MIAs relying on post-hoc data collection
@@
Two concurrent works [85, 68] first developed strategies to create a set of likely members and likely non-members using information about training data sources and the LLM release date. We refer to such data collection strategies as post-hoc, where an attacker aims to infer membership for a released LLM, without access to the randomized training data distribution. In both works, members are collected by sampling documents from the often publicly available datasets widely used for LLM training (e.g. Wikipedia articles [85], books in the public domain, ArXiv papers [68]). Non-members are collected as documents from the same data sources but released after the LLM’s training data cutoff date. Using this collection technique for evaluation, Shi et al. [85] propose a novel MIA scoring function based on the lowest predicted probabilities returned by the LLM. Authors achieved a membership AUC of $0.74$ for LLaMA-30B [93] on the WikiMIA dataset, and an AUC of $0.88$ for GPT-3 [9] on copyright-protected books (BookMIA). Meeus et al. [68] develop a custom membership classifier against OpenLLaMA [32] and achieved MIA AUC of $0.86$ and $0.68$ for books and academic papers, respectively.
@@
Such post-hoc approaches to collect datasets for MIA evaluation are not randomized, raising the question whether they sufficiently approximate randomness in a privacy game and if biases may be introduced during the data collection process. Indeed, doubts have recently been raised on whether this construction of members and non-members constitutes a distribution shift due to the temporal gap in data collection [27, 65], even if efforts to mitigate any distribution shift are employed (e.g. Meeus et al. [68] ensured that the original publication dates of member and non-member books were similarly distributed). If a distribution shift were to be present, the membership classifier might learn the temporal gap in language rather than membership, potentially questioning the validity of MIA results obtained.
IV-B Quantifying a potential distribution shift
@@
Duan et al. [27] was the first to flag that the previously reported high MIA AUC might be due to the distribution shift between members and non-members. They observe a significant change in $7$-gram overlap between members and non-members, and also show how the MIA AUC increases when the temporal gap between members and non-members increases. Shortly afterwards, Maini et al. [65] also stated that post-hoc data collection may lead to a confounding variable, making members and non-members not IID.
@@
We here introduce a bag of words (BoW) classifier to distinguish members from non-members, without taking the target model from the MIA into account. The performance of this classifier reflects what is achievable purely based on the documents, and thus acts as a model-less baseline for the MIA. If this classifier outperforms random guessing in the binary classification task based solely on the dataset’s vocabulary, it may indicate a language shift.
@@
We use a random forest classifier ($500$ trees with maximum depth of $2$ and minimum samples per leaf of $10$) with the count of words ($1$-grams) appearing in at least $5\%$ of the training documents as features. For each dataset, we train the classifier on $80\%$ of all documents, and evaluate it on the remaining $20\%$ (averaged over $5$ runs, with balanced membership). To understand the classifier’s performance, we also identify the words contributing most significantly to the classifier’s decisions.
@@
We instantiate the bag of words classifier on all (available) datasets collected post-hoc and used to evaluate MIAs against LLMs. The WikiMIA and BookMIA datasets, introduced by Shi et al. [85], contain excerpts (of $32$ and $128$ tokens) from Wikipedia ‘events’ and copyright protected books before and after early 2023. Gutenberg (full), introduced by Meeus et al. [68] contains full books released on Project Gutenberg before and after the release date of OpenLLaMA [32], while Gutenberg (mid) contains the same books but with the first and last 10% removed. We also use ArXiv from Meeus et al. [68], containing academic papers before and after the release date of OpenLLaMA [32]. Similarly, ArXivMIA as introduced by Liu et al. [60] contains paper abstracts before and after early 2024. Lastly, we consider StackMIA, as introduced by Ye et al. [104], which contains samples from Stack Exchange before and after early 2023.
TABLE II: AUC (mean±$\pm$±standard deviation) for the bag of words (BoW) classifier across datasets collected post-hoc.
Dataset BoW AUC Predictive words (sample)
@@
Table II shows that the bag of words classifier can distinguish members from non-members significantly better than a random guess (AUC of $0.5$). For some datasets, the classifier reaches near-perfect AUC, i.e. $0.99$ for the WikiMIA dataset, $0.94$ for the BookMIA dataset and $0.97$ for the books from Project Gutenberg (full), strikingly exceeding the MIA AUC reported by attacks utilizing the corresponding target model [85, 68]. For other datasets, the performance is relatively lower, yet remains significantly better (AUC of $\geq 0.6$) than a random guess baseline. This means that, all datasets relying on post-hoc data collection which have been released, suffer from a significant and often strong distribution shift. Moreover, the simplicity of the bag of words classifier suggests that it likely provides a lower bound on classification performance. More advanced classifiers could potentially detect other, subtler distribution shifts.
@@
For WikiMIA [85], most predictive words correspond to years. As the dataset is constructed using text from Wikipedia ‘events’, entries almost always contain dates. Data collected for different time periods could thus be distinguished using the event date mentioned in the excerpts. This is precisely what the bag of words classifier exploits, as its top predictive words are primarily years. Similarly for ArXiv papers used by Meeus et al. [68], years (in citations) carry the most predictive value. For ArXivMIA [60], the bias in years gets eliminated (they use abstracts, which rarely contain dates). Yet, we recover an AUC of $0.60$ with the most predictive words suggesting a change in topics over time (increased use of ‘training’ and ‘github’ in recent work).
@@
In one of the initial MIAs evaluated on data collected post-hoc, Meeus et al. [68] even attempted to control for a potential distribution shift in books by ensuring publication year to be similarly distributed between 1850 and 1910 for members and non-members. Yet, the AUC (Gutenberg (full)) remains near-perfect with the most predictive words relating to template text or formatting. We find that the beginning and end of the books from Project Gutenberg include certain templates (e.g. ‘This ebook was transcribed‘), which have changed over time, allowing the classifier to distinguish members from non-members. Similarly for BookMIA [85], certain formatting styles have changed between members and non-members (e.g. ‘did not‘ rather than ‘didn’t‘).
@@
The distribution shift can be very subtle. We attempt to mitigate the distribution shift due to the template text by excluding the first and last 10% of the books (Gutenberg (mid)). Table II shows, however, that the AUC of the bag of words classifier does decrease from $0.97$, yet remains as high as $0.89$. The set of predictive words shows that books, while originally published in the same time period but released online at a later stage, contain less old formatting and notably, less politically incorrect words such as the n-word.
@@
Overall, these results show that a simple bag-of-words classifier can consistently distinguish between members and non-members across all datasets. This implies that a fundamental assumption in membership inference, i.e. the members and non-members are IID, does not hold and that the post-hoc collection of members and non-members fails to provide a proper privacy game to evaluate MIAs. We are hence unable to exclusively attribute the high MIA AUC reported by studies relying on post-hoc data collection [68, 85, 109, 100, 60, 104, 107, 96, 72, 45] to memorization of the LLM.
IV-C Perpetuating flaws in subsequent research
@@
Since the two initial pieces of work relying on data collected post-hoc [85, 68], we identify at least 10 novel attack methodologies evaluated using datasets collected post-hoc [82, 109, 100, 60, 28, 104, 107, 45, 96], the latest of which proposed (shortly before submission) in September ’24 [96]. These works either use datasets released by Shi et al. [85] (WikiMIA and BookMIA) or new datasets of members and non-members created in a similar, post-hoc manner: ArXivMIA [60], StackMIA [104], Dolma-Book [107] and Wikipedia Cutoff [72] (the latter two of which have not been made available).
@@
With the results of our bag of words classifier (Table II), we find that all (accessible) datasets collected post-hoc suffer from distribution shifts between members and non-members. Therefore, we can not reliably determine whether an MIA achieves a certain performance because it identifies which data points are memorized by the target model or simply due to its ability to detect the distribution shift. Likewise, any comparison of MIA methodologies evaluated with these datasets becomes unreliable.
@@
The problem also extends beyond the development of new MIAs, with subsequent work using datasets collected post-hoc to evaluate the effectiveness of watermarking techniques [48, 99] as a privacy-preserving technology [78] or the success of machine unlearning [102]. The use of flawed MIA evaluation setups in these contexts is highly concerning. Indeed, it could lead to false conclusions about the effectiveness of privacy-preserving technologies, which could have far-reaching implications for LLM development and deployment.
V Randomization for MIAs against LLMs
@@
V-A Why is it different than for MIAs against ML models?
@@
MIAs have been extensively applied to various domains of ML, including image classification [87, 13, 83, 95, 59], synthetic data generation [35, 90, 67] and natural language processing [84, 36, 92, 88], all evaluated in randomized privacy games. In the fixed record privacy game [103], multiple ML models are trained with the target record either included or excluded, commonly referred to as shadow modeling [61, 62, 90, 41, 11, 81, 105]. For instance, Shokri et al.[87] used up to $100$ shadow models for image classification models, while Song et al.[89] trained $10$ shadow models to evaluate privacy in smaller language models (RNNs). However, replicating fixed record privacy games for LLMs is impractical due to the extremely high costs of pretraining.
@@
As a result, privacy games to evaluate MIAs against LLMs likely rely on the fixed model scenario, in which a single ML model is trained with randomly sampled target records from the same distribution, either included or excluded. Yet, the practical implementation of fixed model privacy games for LLMs comes with its limitations, which we will address below.
V-B Unit of interest to infer membership
@@
In traditional MIAs against ML models, the unit of interest to infer membership is naturally defined (e.g. single record, image). For text, most prior work considered sequences of a fixed, often truncated length as the unit of interest. For instance, Carlini et al. [15] use sequences of $256$ tokens, while Shi et al. [85] and subsequent work [109, 100, 60, 104, 107, 96, 72, 45] consider random excerpts of $32$ to $256$ tokens. When injecting randomized sequences, prior work has considered sequences up to $80$ [99] or $100$ [69] tokens.
@@
In some cases, however, the goal of applying MIAs against LLMs is to infer membership of larger pieces of text, from the level of a document [68, 85] to an entire dataset [65]. For instance, in the context of copyright, the goal is to infer whether the entire content (e.g. news articles, book) was used rather than certain excerpts. This has motivated prior work [68, 85, 65] to instantiate document- or even dataset-level membership inference, aiming to aggregate membership signal to larger pieces of text.
@@
Formally, a document-level MIA aims to determine whether a given document $d_{t}=(t_{1},t_{2},...,t_{N_{d_{i}}})$ containing $N_{d_{t}}$ tokens, was included in the training dataset $\mathcal{D}$ for target LLM $\mathcal{M}$. The MIA defines a scoring function $f_{\text{doc}}(\mathcal{M},d_{t})$ such that $\hat{y}_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }f_{\text{doc}}(\mathcal{M},d_{t})\geq\tau_{\text{doc}}\text{, and }0\text{ otherwise}$. The true membership label is $y_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }d_{i}\in\mathcal{D}\text{, and }0\text{ otherwise}$. For simplicity, we argue that the task of dataset inference as proposed by Maini et al. [65] is a generalization of document-level MIAs.
@@
Document-level MIAs differ from sequence-level MIAs against LLMs, as aggregating meaningful signal across significantly more tokens might be challenging. Multiple ways have been proposed to compute $f_{\text{doc}}(\mathcal{M},d_{t})$. Meeus et al. [68] iterates through the entire document $d_{t}$, queries the LLM on truncated chunks, and aggregates all token-level probabilities to multiple document-level features as input for a membership classifier. Shi et al. [85] proposes to sample $k$ random excerpts, i.e. sequences $s_{tj}$, from document $d_{t}$ to compute the sequence-level scoring function $f_{\text{seq}}(\mathcal{M},s_{tj})$ for all $j=1,\ldots k$. This document-level scoring function is then computed as an average binary prediction $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$, where $\tau_{\text{seq}}$ is determined to maximize accuracy on the sequence-level MIA. Lastly, Maini et al. [65] aggregates values of the sequence-level scoring function to a document-level statistical test.
V-C Overlap between members and non-members